# Transforming Role Classification in Scientific Teams Using LLMs and Advanced Predictive Analytics

### v1

@@ 1 Introduction
We present a comprehensive framework that utilizes such LLMs to classify roles within scientific teams. Our methodology includes a few-shot prompt to accurately categorize author roles. Furthermore, after the LLM inference step, we aim at feature engineering: the framework incorporates ten extracted features reflecting each author’s contributions and characteristics, such as contribution to references, contribution to topics, probability of leading, probability of managing correspondence, career age, citation count, unique topics, total publications, citation impact per year, and institutional diversity, by using a dataset derived from the OpenAlex database, which provides extensive metadata on academic publications, including citation counts, author affiliations, research topics, and publication history [8]. We further train a deep learning model using 10 extracted features to classify author roles with high accuracy. This approach not only complements, but also improves upon traditional methods, providing a more detailed and accurate classification of author roles. By leveraging the advanced capabilities of LLMs, our research aims to transform the landscape of scientific team role classification and provide a more nuanced understanding of team structures and collaboration dynamics.
@@ 3 Dataset
We first utilize self-reported data collected from papers published in prominent journals such as PNAS, Nature, Science, and PLOS One from 2003 to 2020, which was opened by Xu et al [2]. These four journals were selected due to their high impact and influence in the scientific community, ensuring high quality and significant research data. In addition, recent efforts to make scientific research fairer and more transparent have led these leading journals to require detailed reporting of each author’s specific contributions. This requirement helps ensure that everyone involved in the research receives proper credit, making these journals ideal for our analysis. This selection provides a representative dataset that supports future work to understand and improve the dynamics of scientific collaboration.
@@
4 Role Classification
In this stage, we focus on classifying and categorizing roles within scientific teams using self-reported data. Our methodology involves the use of open-source instructional LLMs and the GPT-4 model, augmented with few-shot prompting. Through this process, we aim to achieve a detailed categorization of scientific roles, thereby improving our understanding of team dynamics in scientific research environments.
4.1 Initial Role Classification Via LLMs
@@
Figure 1: Illustration of the workflow involving data sampling, preprocessing, contribution assigning, and classification task.
@@
4.2 Methodology for Role Classification using LLMs
@@
To facilitate this, we created specific prompts for the models, guiding them to categorize and classify research roles based on the activities associated with each role, as illustrated in Figure 2:
@@
4.3 Model Configuration and Results
@@ 4.4 Comparison of LLMs with Traditional Models
Table 2: Performance of BERT and RoBERTa models.
@@
5 Role Classification Predictive Model
5.1 Dataset and Feature Engineering
After the initial role classification using LLMs, to build our predictive model, we first constructed a robust dataset derived from the OpenAlex database, which provides extensive bibliometric and metadata on academic publications. We then moved on to building a predictive role classification model using feature engineering.
@@ 10 Collaboration Diversity Metrics Institutional Diversity
5.2 Normalization
Normalizing the dataset was a critical step in preparing it for predictive modeling. We applied a min-max scaling to each feature to ensure that they contributed equally to the model, without any single feature disproportionately influencing the results due to its size. The formula used for min-max normalization is given by:
@@ 5.2 Normalization
5.3 Data Splitting and Modeling
To ensure a robust evaluation of the prediction models, we expanded our dataset by further inferring the roles of 10,000 distinct authors using the classification framework developed in the first phase, resulting in a total dataset of about 15,000 data points. Originally, the dataset included three classes: Leadership, Direct Support, and Indirect Support. For the purpose of this binary classification task, we merged the Direct Support and Indirect Support roles into a single “Support” category, while maintaining “Leadership” as a distinct category.
This expanded dataset included detailed role classifications based on each author’s contributions. The dataset was then divided into training and test sets using stratified sampling to maintain the distribution of the target variable. A ratio of 0.8 to 0.2 was used for the split, stratifying by two types of roles classified as leadership and non-leadership types.
5.4 Model Training and Results
@@ 6 Conclusions
In addition, this methodology allows us to further calculate the L-Ratio, which quantifies the proportion of leadership roles within research teams. In future work, by integrating the L-Ratio with other variables, we can explore the relationship between leadership dynamics and various aspects of academic performance. For example, using the L-Ratio to analyze the long-term performance of junior team members can provide insights into their career trajectories and contributions over time. This analysis aims to highlight the importance of leadership dynamics in scientific collaboration, providing valuable information on how junior researchers’ roles and impacts evolve within academic teams.
Despite these promising results, our study has limitations that suggest directions for future research. While advanced LLMs such as GPT-4 demonstrate superior performance in author role classification, their proprietary nature and computational requirements may limit accessibility for some researchers. Additionally, our reliance on data derived from the OpenAlex database may not capture the full diversity of author contributions across all scientific disciplines and publication venues. To address these limitations, future work could involve exploring zero-shot inference with LLMs to obtain more natural and pattern-rich results. Furthermore, analyzing the patterns and insights derived from zero-shot inference could enhance our understanding of author roles and contributions, particularly in handling complex or ambiguous cases that challenge traditional rule-based approaches.

### v2

@@ 1 Introduction
We present a comprehensive framework that utilizes such LLMs to classify roles within scientific teams. Our methodology includes a few-shot prompt to accurately categorize author roles. Furthermore, after the LLM inference step, we aim at feature engineering: the framework incorporates ten extracted features reflecting each author’s contributions and characteristics, such as contribution to references, contribution to topics, probability of leading, probability of managing correspondence, career age, citation count, unique topics, total publications, citation impact per year, and institutional diversity, by using a dataset derived from the OpenAlex database, which provides extensive metadata on academic publications, including citation counts, author affiliations, research topics, and publication history [8]. We further train a deep learning model using 10 extracted features to classify author roles with high accuracy. Moreover, we conducted SHAP (SHapley Additive exPlanations) [9] analysis to investigate the importance of these features in the prediction tasks, providing valuable insights into their contributions. This approach not only complements, but also improves upon traditional methods, providing a more detailed and accurate classification of author roles. By leveraging the advanced capabilities of LLMs, our research aims to transform the landscape of scientific team role classification and provide a more nuanced understanding of team structures and collaboration dynamics.
@@ 3 Dataset
We first utilize self-reported data collected from papers published in prominent journals such as PNAS, Nature, Science, and PLoS One from 2003 to 2020, which was opened by Xu et al [2]. These four journals were selected due to their high impact and influence in the scientific community, ensuring high quality and significant research data. In addition, recent efforts to make scientific research fairer and more transparent have led these leading journals to require detailed reporting of each author’s specific contributions. This requirement helps ensure that everyone involved in the research receives proper credit, making these journals ideal for our analysis. Specifically, PLoS One was included as it mandates comprehensive author contribution statements similar to the other journals, providing rich and structured data for analysis. Additionally, its broad scope across various scientific disciplines enables a more diverse representation of research team structures. To ensure consistency and reduce selection bias, we followed the methodology of Xu et al. [2] and stratified our sampling across these four journals. This selection provides a representative dataset that supports future work to understand and improve the dynamics of scientific collaboration.
@@
4 LLM-Based Role Classification: Methods and Evaluation
In this stage, we focus on classifying and categorizing roles within scientific teams using self-reported data. Our methodology involves the use of open-source instructional LLMs and the GPT-4 model, augmented with few-shot prompting. Through this process, we aim to achieve a detailed categorization of scientific roles, thereby improving our understanding of team dynamics in scientific research environments. This classification serves as a crucial foundation for subsequent predictive modeling. While LLMs provide high-quality classifications, their computational demands make them impractical for large-scale applications. To address this, the LLM-generated labels will be leveraged to train a predictive model capable of efficiently scaling to extensive datasets. Section 5 details this predictive modeling process, which extends our role classification framework to broader scientific collaborations.
4.1 Overview of LLM-Based Role Classification Tasks
@@
Figure 1: Illustration of the workflow involving data sampling, preprocessing, contribution assigning, and classification task for the LLM-based role classification.
@@
4.2 Prompt Engineering for Role Classification
@@
To facilitate this, we created specific prompts for the models, guiding them to categorize and classify research roles based on the activities associated with each role, as illustrated in Figure 2. We also explored zero-shot prompting, where the model relies solely on its pre-trained knowledge without additional examples. However, we observed that the model’s predictions in zero-shot mode were less stable, often producing inconsistent or ambiguous role assignments. Without explicit demonstrations, LLMs sometimes misclassified contributions, particularly when role descriptions were subtle or overlapping. In contrast, our few-shot approach, which provided targeted examples, significantly improved classification accuracy by reducing ambiguity and ensuring better role differentiation.
To construct effective prompts, we defined a structured set of role categories and associated key responsibilities, ensuring clarity in classification. These role definitions provided the necessary context for LLMs to accurately assign contributions, as detailed below:
@@
4.3 LLM Model Configuration and Performance Comparison
@@ 4.4 Comparison of LLMs with Traditional Models
Table 2: Performance of BERT and RoBERTa Models on Role Classificaiton.
@@
5 Scalable Predictive Modeling for Author Role Classification
To enable large-scale and efficient analysis of author roles in scientific collaborations, we address the computational challenges associated with directly applying GPT-4 for large-scale predictions. While GPT-4 demonstrates exceptional performance in classifying author roles, its inference is computationally expensive and impractical for analyzing millions of authors. To overcome this challenge, we utilize labels generated from GPT-4 to train a dense neural network, which enables efficient classification of new data with minimal computational overhead.
5.1 Dataset Construction and Feature Engineering
To build our predictive model, we first constructed a robust dataset derived from the OpenAlex database, which provides extensive bibliometric and metadata on academic publications. We then moved on to building a predictive role classification model using feature engineering.
@@ 10 Collaboration Diversity Metrics Institutional Diversity
5.2 Data Splitting and Normalization
To ensure a robust evaluation of the prediction models, we expanded our dataset by further inferring the roles of 10,000 distinct authors using the classification framework developed in the first phase, resulting in a total dataset of about 15,000 data points. Originally, the dataset included three classes: Leadership, Direct Support, and Indirect Support. For the purpose of this binary classification task, we merged the Direct Support and Indirect Support roles into a single “Support” category, while maintaining “Leadership” as a distinct category. This expanded dataset included detailed role classifications based on each author’s contributions. The dataset was then divided into training and test sets using stratified sampling to maintain the distribution of the target variable. A ratio of 0.8 to 0.2 was used for the split, stratifying by two types of roles classified as leadership and non-leadership types.
Additionally, Normalizing the dataset was a critical step in preparing it for predictive modeling. We applied a min-max scaling to each feature to ensure that they contributed equally to the model, without any single feature disproportionately influencing the results due to its size. The formula used for min-max normalization is given by:
@@ 5.2 Data Splitting and Normalization
5.3 Predictive Model Training and Performance
@@
5.4 Interpretability Analysis Using SHAP
To further validate feature importance in our classification task, we apply SHAP (SHapley Additive exPlanations) [9, 44] to investigate how each predictive feature contributes to the model’s decision-making process. Specifically, we utilize Gradient SHAP, which efficiently estimates feature attributions for deep learning models using gradient-based sampling. the SHAP value for a feature $i$ is calculated as:
$\phi_{i}=\sum_{S\subseteq F\\ i\notin S}\frac{|S|!(M-|S|-1)!}{M!}\left[f(S\cup\{i\})-f(S)\right]$
where:
  • $F$ is the set of all features,
  • $S$ is a subset of features excluding $i$,
  • $f(S)$ is the model’s prediction using only the features in $S$,
  • $\phi_{i}$ is the SHAP value for feature $i$.
Figure 5: SHAP summary plot showing feature importance and the directional impact of each feature on the model’s predictions.
As shown in Figure 5, Probability of Leading Correspondence and Probability of Leading emerge as the most influential features in distinguishing leadership roles. These features likely capture the responsibility and visibility of an author within a research project, suggesting that consistent first or corresponding authorship is a strong indicator of leadership roles. Their high SHAP values further confirm their direct impact on classification predictions, reinforcing their centrality in the model’s decision-making process.
Subsequently, Contribution to References and Contribution to Topics also hold substantial importance, showcasing their role in capturing intellectual contributions. A higher contribution to references indicates an author’s influence on shaping discussions within a paper, while topic contributions suggest an author’s role in guiding the thematic direction of research. The positive SHAP impact of these features suggests that intellectual involvement plays a critical role in distinguishing different author roles.
Additionally, the newly proposed features in this paper, including: Citation Impact per Year and Institutional Diversity, provide meaningful contributions. Citation Impact per Year accounts for the sustained influence of an author’s research over time, making it a valuable indicator of long-term academic impact. Its moderate importance in the SHAP analysis suggests that while citations are significant, they are not the sole determinant of leadership roles. Institutional Diversity, on the other hand, captures the extent of an author’s research collaborations across different institutions. A higher value for this feature suggests exposure to diverse academic environments, which may contribute to broader collaborative roles rather than direct leadership.
@@ 6 Conclusions
In addition, this methodology allows us to further calculate the L-Ratio, which quantifies the proportion of leadership roles within research teams. In future work, by integrating the L-Ratio with other variables, we can explore the relationship between leadership dynamics and various aspects of academic performance. For instance, using the L-Ratio to analyze the long-term performance of junior team members can provide insights into their career trajectories and contributions over time. This analysis aims to highlight the importance of leadership dynamics in scientific collaboration, providing valuable information on how junior researchers’ roles and impacts evolve within academic teams.
@@
7 Discussion: Technical and Practical Implications
Our study introduces a technical framework that leverages few-shot prompting with LLMs and a scalable neural network to generate high-quality labels without extensive training data. The use of SHAP for interpretability highlights the key factors—such as the probability of leading—that define leadership roles. Practically, our approach supports a deeper understanding of scientific collaboration. Moreover, the scalable predictive model paves the way for automated monitoring of author contributions across large datasets.
Despite these promising results, our study has limitations that suggest directions for future research. While advanced LLMs such as GPT-4 demonstrate superior performance in author role classification, their proprietary nature and computational requirements may limit accessibility for some researchers. Additionally, our reliance on data derived from the OpenAlex database may not capture the full diversity of author contributions across all scientific discipline and publication venues. To address these limitations, future work could involve exploring zero-shot inference with LLMs to obtain more natural and pattern-rich results. Furthermore, analyzing the patterns and insights derived from zero-shot inference could enhance our understanding of author roles and contributions, particularly in handling complex or ambiguous cases that challenge traditional rule-based approaches.